Determine the utilization and the efficiency for each of these situations:

a. A loan processing operation that processes an average of 5 loans per day. The operation has a design capacity of 11 loans per day and an effective capacity of 9 loans per day. (Round your answer to 1 decimal place. Omit the "%" sign in your response.)

Utilization %
Efficiency %


b. A furnace repair team that services an average of 4 furnaces a day if the design capacity is 8 furnaces a day and the effective capacity is 6 furnaces a day. (Round your answer to 1 decimal place. Omit the "%" sign in your response.)

Utilization %
Efficiency %


c. Would you say that systems that have higher efficiency ratios than other systems will always have higher utilization ratios than those other systems?

This is not necessarily  (Click to select)  false  true  . If the design capacity is relatively  (Click to select)  low  high  , the utilization could be  (Click to select)  high  low  even though the efficiency was  (Click to select)  low  high  .
Answer #1

Here is the answer:

Utilization %= Output/Design Capacity

Efficiency %= Effective capacity/Design Capacity

Hence,

A. U= 5/11 = 45.5, E= 9/11 = 81.8

B. U= 4/8= 50, E= 6/8= 75

C. This is not necessarily TRUE. If the design capacity is relatively HIGH, the utilization could be LOW even though the efficiency was HIGH.